Brenton D. Walker

1130 Erin St. Apt. 1 Madison, WI 53715 (608) 286-0416

email: bdwalker@students.wisc.edu

Objective Challenging, computer related position. Education

• 6/95 Diploma, John Marshall High School, Milwaukee, WI – 3.7 GPA

– National Honor Society 1994, 1995

• University of Wisconsin–Milwaukee – 6/94 – 6/95 PSEO High School Special Student

– 4.0 GPA

• University of Wisconsin–Madison – 5/95 Admitted into Honors Program

– Major in Math, Computer Science, and Chemistry

– 3.8 Cumulative GPA

– Planned Graduation Date: December 2000

Skills and Abilities • Programming: Basic, C++, Fortran, HTML, Java, L^{A}TEX, Lisp, Pascal, Perl

• Operating Systems: MacOS, Linux, Unix, VMS, DOS, Windows 3.1/95/NT, Cisco IOS Experience

• 1/95 – 6/95 Organizer, Marshall High School Computer Networking Club – Taught students and teachers about how to use the Internet.

• 2/96 – 6/96 InfoLab Consultant, Social Sciences Microlab

• 2/96 – 6/96 Computer Operator, Social Science Computing Coop – Monitored Unix, VMS, and NT systems

– Mounted tapes

– Performed various syetem administration tasks and wrote Perl programs

• 6/96 – 4/99 Network Assistant, Social Science Computing Coop – Configured, installed, and supported network equipment

– Troubleshoot problems involving Ethernet, FDDI, ATM, IP, RS232

– Wrote network management applications

– Installed and maintained data wiring

• 4/97 – 6/98 Webmaster, Social Sciences Microlab

• 4/99 – present Network Manager, Madison Property Management Inc and Madison Telecom Inc.

– Set up and maintain NT3, NT4 servers and Win95/98 client machines

– Set up and maintain Linux servers, DNS records, Sendmail, Apache – Configure and maintain IP firewlls

– Travel to customer sites to install new software and hardware

Honors and Activities • American Chemical Society Award for Excellence in Physical Chemistry 5/1999 • Elected to Phi Beta Kappa 2/1999

• Member, Brazen Dropouts Cycling Club, 7/1997 – present

• Member, UW Cycling Club, 1/1997 – present

• Marshall High School Math Award 1994, 1995

• Air Force Recruiting Service Math and Science Award 1995

• Bausch & Lomb Honorary Science Award 1994 Interests Cycling and bike repair, writing small math-intensive programs, Internet surfing, and

reading. References Available upon request.